OpenStudy (anonymous):

HELP please ??? kathy wants to factor x^2 - 12x + 36 by grouping; however, Tom says it is a special product and can factor a different way. Using complete sentences,explain and demonstrate how both methods will be result in the same factor

OpenStudy (gorv):

first x^2-6x-6x+36=x(x-6)-6(x-6)=(x-6)*(x-6) second it is a perfect square x^2-2*6*+6^2=(x-6)^2

OpenStudy (gorv):

@savannaxx_ got it???

OpenStudy (anonymous):

you could use the method by factorisation where it will be x^2-6x-6x+36 which gives (x-6)(x-6) as factors

OpenStudy (anonymous):

its true gorv got it

OpenStudy (anonymous):

thanks you guys :) @nelotsak @gorv